OpenStudy (itsbribro):

What is the value of x? Round your answer to the nearest tenth if necessary.

OpenStudy (liamschumm):

So, do you know what the Pythagorean Theorem states?

OpenStudy (anonymous):

Use the pythagorean theorem so that a^2+b^2=c^2

OpenStudy (anonymous):

the hypotenuse is always C

OpenStudy (liamschumm):

Yes, j.

OpenStudy (liamschumm):

Itsbribro, do you get it?

OpenStudy (liamschumm):

Hello?

OpenStudy (itsbribro):

yes i tried that i dont understand

OpenStudy (liamschumm):

Let us apply this.

OpenStudy (liamschumm):

6^2 + 4^2 = x^2

OpenStudy (liamschumm):

Calculate the left hand side of that for me, please?

OpenStudy (itsbribro):

6^2= 36 and 4^2=16

OpenStudy (liamschumm):

So, add them.

OpenStudy (itsbribro):

sorry internet sucks

OpenStudy (liamschumm):

Nah, it is fine.

OpenStudy (itsbribro):

36+16=52

OpenStudy (liamschumm):

Great!

OpenStudy (liamschumm):

So, get your calculator and take the square root of 52.

OpenStudy (itsbribro):

7.2?

OpenStudy (liamschumm):

Yep.

OpenStudy (liamschumm):

Great work! Give me a medal if I helped you out.

OpenStudy (liamschumm):

Have a great day!

OpenStudy (itsbribro):

you too thank you so much